Tanh x=3/5 Tanh 3x=? a)63/65 b)61/65 c)65/61 d)65/63 Plz explain..

Tanh x=3/5


Tanh 3x=?


a)63/65


b)61/65


c)65/61


d)65/63


Plz explain..

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

tanh x=(e^x-e^-x)/(e^x+e^-x)=(e^2x -1)/(e^x +1)=3/5,5(e^2x -1)=3(e^2x +1), 2e^2x =8, e^2x =4, e^x=2,tanh 3x=

(8-(1/8))/(8+(1/8))=63/65